### v1

# From RNNs to Foundation Models: An Empirical Study on Commercial Building Energy Consumption

Abstract
Accurate short-term energy consumption forecasting for commercial buildings is crucial for smart grid operations. While smart meters and deep learning models enable forecasting using past data from multiple buildings, data heterogeneity from diverse buildings can reduce model performance. The impact of increasing dataset heterogeneity in time series forecasting, while keeping size and model constant, is understudied. We tackle this issue using the ComStock dataset, which provides synthetic energy consumption data for U.S. commercial buildings. Two curated subsets, identical in size and region but differing in building type diversity, are used to assess the performance of various time series forecasting models, including fine-tuned open-source foundation models (FMs). The results show that dataset heterogeneity and model architecture have a greater impact on post-training forecasting performance than the parameter count. Moreover, despite the higher computational cost, fine-tuned FMs demonstrate competitive performance compared to base models trained from scratch.
1 Introduction
In smart grids, accurate forecasting of future load consumption across different timescales is critical for both operational tasks, such as generator scheduling, and long-term planning, such as capacity expansion. A key area of focus is short-term load forecasting (STLF), which predicts energy consumption over timeframes from an hour to a day, at the level of individual buildings (Espinoza et al., 2007). The canonical pipeline for STLF involves collecting smart meter data of multiple buildings into a single dataset, training a single model on the same, and validating it on data split along the time axis (Kong et al., 2019). Pooling large datasets is particularly crucial for training large models, including foundation models, as it enables the models to learn diverse patterns and complex relationships from a wide variety of buildings. This global approach of training on pooled datasets requires forecasting models to effectively manage the heterogeneity that arises from differences in building data (Hewamalage et al., 2022). The heterogeneity may manifest in scale, variance, seasonality, or trend of different time series.
We use the ComStock dataset (Parker et al., 2023) to study the effect of heterogeneity on various STLF models. ComStock is a synthetic dataset containing energy consumption of more than 350,000 U.S. commercial buildings over one year, alongside additional features. These features include building type, allowing us to control heterogeneity in curated ComStock subsets by varying building type diversity. We generate two such datasets, with one exhibiting greater heterogeneity. In this context, we evaluate multiple time series forecasting models from the literature, broadly categorized into two types: base models and foundation models (FMs). Base models are trained from scratch on a given dataset, while FMs are pre-trained on a large corpora, and only need to be fine-tuned for the given downstream application. We sample across a wide cross-section of available base models: recurrent neural networks (RNNs) (LSTM (Hochreiter and Schmidhuber, 1997), LSTNet (Lai et al., 2018)), transformer-based architectures (Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021)), decomposition-based architectures (Autoformer (Wu et al., 2021)), and patch-based architectures (TimesNet (Wu et al., 2023)). For foundation models, we select TimesFM (Das et al., 2024), which handles multivariate inputs and is tested in both zero-shot (ZS) and fine-tuned (FT) modes. This setup enables us to assess the benefits of pre-training alongside the impact of heterogeneity.
This paper makes three key contributions: (i) the introduction of two curated building energy consumption datasets from ComStock, which are of the same size but vary in heterogeneity; (ii) a performance comparison of various base models on these datasets; and (iii) an evaluation of foundation models in both zero-shot and fine-tuned modes, compared to base models.
2 ComStock Dataset
Figure 1: (a) Distribution of building type in IL-HET. (b) Boxplot of the spread of load values for each building type in IL-HET. (c) Load shapes from two buildings each from IL-HET and IL-HOM.
ComStock, developed by the National Renewable Energy Laboratory and funded by the U.S. Department of Energy, is a detailed model of the U.S. commercial building stock. It uses statistical sampling methods and advanced energy simulations via OpenStudio (Guglielmetti et al., 2011) to estimate energy consumption across more than 350,000 unique building energy models. These models encompases various building characteristics, including building type, construction year, HVAC system, and occupancy schedules, using 15-minute resolution data for one year. The dataset provides detailed outputs like energy consumption across different end uses and critical metrics, resulting in approximately 10 TB of data per release.
For our study, we distill the ComStock data into two curated datasets from buildings in Illinois, termed IL-HET (heterogeneous) and IL-HOM (homogeneous). Each dataset contains 592 buildings for the actual meteorological year 2018. IL-HET includes all 14 available building types in ComStock (see Figure 2(a)), maintaining the overall dataset distribution, thus introducing heterogeneity. In contrast, IL-HOM comprises only buildings of type Warehouse, providing a more homogeneous set due to correlations among similar building loads. The heterogeneity in IL-HET is illustrated through its temporal spread in Figure 2(b), while IL-HOM demonstrates a more uniform pattern (Appendix A).
Apart from the load in kWh, we incorporate several features in the dataset. We incorporate temporal information using two categorical features that represent the index of the 15-minute interval of the day and the day of the week. As mentioned earlier, we include the actual meteorological year 2018 weather features by the way of temperature and wind speed. Furthermore, we include three building-specific static features viz floor space, wall area, and window area. This feature selection is broader than the one in conventional energy datasets (see Appendix A), representative of the typical data used by utility companies (Jawed, 2018), and is validated via correlation analysis on a separate selection of buildings not included in either IL-HET or IL-HOM (see Appendix B).
3 Experiment Design
A pointwise time series forecasting problem can be generally represented as
$\hat{\mathbf{y}}_{t+T}=f_{\boldsymbol{\theta}}(\mathbf{y}_{[t-L+1:t]},\mathbf{u}_{[t-L+1:t+T]},\mathbf{x}_{[t-L+1:t]},\mathbf{s}),$
where $\mathbf{y}_{t}\in\mathbb{R}^{n}$ is the building energy consumption, $\mathbf{x}\in\mathbb{R}^{m}$ includes the two weather features, $\mathbf{u}_{t}$ denotes two time induces, and $\mathbf{s}$ includes the three static building characteristics. Furthermore, $T$ and $L$ are the ‘lookahead’ and ‘lookback’ durations, respectively. The forecast model $f(\cdot)$ that is parameterized by the vector $\boldsymbol{\theta}\in\mathbb{R}^{h}$ can be trained in a supervised fashion, followed by evaluating its performance on unseen test sets. We use two metrics: mean squared error (MSE) and mean absolute error (MAE), which given time indices of interest $t\in\mathcal{T}$, are defined as $\text{MSE}:=\frac{1}{n|\mathcal{T}|}\sum_{t\in\mathcal{T}}\left\lVert\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\right\rVert_{2}^{2}$ and $\text{MAE}:=\frac{1}{n|\mathcal{T}|}\sum_{t\in\mathcal{T}}\left\lVert\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\right\rVert_{1}$, respectively. While the models are trained using the MSE loss, both MSE and MAE metrics are recorded over the test sets.
The dataset is divided into train, validation, and test sets in a ratio of $0.8:0.1:0.1$ along the time axis for each building. Each set is $z$-normalized using statistics from the train set. Data from all buildings are then concatenated to form unified datasets. Our choice of models (cf. Section 1) aims to provide a comprehensive evaluation of various time series models from the literature. Detailed descriptions of the model categories are provided in Appendix C.
For training the base models and fine-tuning TimesFM, we use the Adam optimizer with default parameters $\beta_{1}=0.9,\beta_{2}=0.999$. Learning rates are tuned per model from the set of $\{10^{-3},5\times 10^{-4},10^{-4},5\times 10^{-5},10^{-5},5\times 10^{-6}\}$, except for Informer and TimesFM. Informer underwent additional tuning on the set $\{10^{-6},5\times 10^{-7},10^{-7}\}$, while TimesFM was also tuned on $10^{-2}$. Base models are trained with large batches of size 1024, and training stops after 20 epochs or if there is no improvement in the MSE metric on the validation set for 5 consecutive epochs.
For fine-tuning TimesFM, we use a smaller batch size of 512 due to its larger memory requirements compared to the base models. The fine-tuning process is performed for 2 epochs, taking a similar amount of time to train each base model for 20 epochs. The zero-shot experiments with TimesFM are performed using the off-the-shelf pre-trained timesfm-1.0-200m (https://huggingface.co/google/timesfm-1.0-200m) without any modification. A detailed description of the model hyperparameters, learning rates, and termination epochs is provided in Appendix D.
4 Experiment Results
The performance of the trained models on the test set is summarized in Table 1, and predictions on the test set are visualized in Appendix E. Immediately, it is obvious that TimesNet provides the best performance for all lookahead values across both the datasets. This can be attributed to the patch-based architecture of TimesNet, which enables it to utilize the information present in features effectively. The superior performance of LSTM and LSTNet as compared to Transformer-based architectures is also notable, and we hypothesize that recurrent architectures of the former utilize feature information more efficiently than the latter’s tokenization and attention schemes. Furthermore, we clearly see the impact of heterogeneity—all experiments with IL-HOM have lower metrics compared to those with IL-HET except for TimesNet and Autoformer—and even here, the difference is minute. This validates our hypothesis that heterogeneous datasets are more difficult to learn than homogeneous datasets, while demonstrating that architectures such as TimesNet are robust to increasing heterogeneity.
Comparing the base models to FMs, we see that TimesFM when fine-tuned offers performance close to the best base models. This demonstrates the inherent benefit of the vast corpora of information absorbed by the model when it comes to downstream forecasting tasks. On the other hand, the zero-shot performance of TimesFM is inadequate compared to all base models, highlighting the importance of fine-tuning to unlock FMs’ performance. We remark that although we fine-tuned the model for only 2 epochs in this study, further fine-tuning with more epochs could potentially improve accuracy. However, this would require significantly more time and computing resources, as compared to the base models.
We remark that some of the base models terminated much earlier than 20 epochs due to the patience criterion. Specifically, several models terminated training in as few as 1, 2, 5, or 7 epochs; see Appendix D. This early termination highlights the efficiency of these base models in adapting to the datasets, contrasting with the longer fine-tuning process required by TimesFM.
Finally, we analyze the importance of base model size on performance in Figure 2. This not only visually reinforces the impact of increasing heterogeneity on model performance, but also highlights the point that model architecture is more important than brute-forcing model size. Evidently, the trend that relatively small models viz. LSTM, LSTNet, and TimesNet can outperform bigger models highlights the importance of efficiently learning temporal relations and feature-dependence on forecasts, rather than increasing parameter sizes.
Table 1: Experiment results for different models tested on the two datasets.
IL-HET
$(L,T)$ LSTM LSTNet Transformer Autoformer
MSE MAE MSE MAE MSE MAE MSE MAE
$(96,4)$ 0.1731 0.1517 0.0571 0.1095 0.2725 0.2850 0.1669 0.2858
$(96,48)$ 0.5276 0.3534 0.1463 0.1732 0.4205 0.3443 0.1635 0.2690
$(96,96)$ 0.2154 0.1772 0.0998 0.1420 0.3533 0.3826 0.3065 0.4101
Informer TimesNet TimesFM (ZS) TimesFM (FT)
MSE MAE MSE MAE MSE MAE MSE MAE
$(96,4)$ 0.4027 0.4758 0.0340 0.0637 0.5503 0.3093 0.0614 0.0876
$(96,48)$ 0.5053 0.6004 0.0773 0.0938 0.7269 0.3405 0.1021 0.1322
$(96,96)$ 0.5480 0.5861 0.0726 0.0954 0.7583 0.3559 0.1319 0.1462
IL-HOM
$(L,T)$ LSTM LSTNet Transformer Autoformer
MSE MAE MSE MAE MSE MAE MSE MAE
$(96,4)$ 0.0830 0.1405 0.0542 0.1243 0.1053 0.1765 0.1751 0.2405
$(96,48)$ 0.2619 0.2885 0.1216 0.1918 0.1748 0.2390 0.2123 0.2572
$(96,96)$ 0.1458 0.2008 0.1180 0.1866 0.1601 0.2288 0.2463 0.2802
Informer TimesNet TimesFM (ZS) TimesFM (FT)
MSE MAE MSE MAE MSE MAE MSE MAE
$(96,4)$ 0.0801 0.1598 0.0285 0.0783 0.7463 0.4837 0.0668 0.1047
$(96,48)$ 0.1220 0.1871 0.0960 0.1459 0.7394 0.4794 0.1060 0.1368
$(96,96)$ 0.1387 0.1983 0.1017 0.1534 0.7392 0.4798 0.1286 0.1573
Figure 2: Test set performance of base models for $L=96$ and (a) $T=4$, (b) $T=48$, (c) $T=96$.
5 Concluding Remarks
In this paper, we tested various base models, as well as a fine-tuned and zero-shot FM on two datasets, one more heterogeneous than the other. The results demonstrate the impact of heterogeneity on model performance, the capability of models like TimesNet to be robust against heterogeneity, the advantage of fine-tuned FMs, and their disadvantage in zero-shot mode. These observations serve towards grounding the problem of heterogeneity in time series models and serve as important stepping stones towards domains such as federated learning, where data heterogeneity is inherent to the domain.
Acknowledgments and Disclosure of Funding
This material is based upon work supported by the U.S. Department of Energy, Office of Science, under contract number DE-AC02-06CH11357. This research used resources of the Argonne Leadership Computing Facility at Argonne National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under contract DE-AC02-06CH11357. We also gratefully acknowledge the computing resources provided on Swing, a high-performance computing cluster operated by the Laboratory Computing Resource Center at Argonne National Laboratory.
Appendix A Analysis of IL-HET and IL-HOM
Table 2: Statistics for loads in IL-HOM and IL-HET.
Dataset Mean (loads) Standard deviation (loads)
IL-HOM 17.4307 24.3129
IL-HET 21.8440 51.3635
Both IL-HET and IL-HOM contain load values of 592 buildings, along with 7 other features. Since the data is recorded at a granularity of $15$ minutes for one year, the total number of data points are 35060. Thus, each of the datasets contain a total of 166,044,160 observations. Considering the fact that both datasets contain the same number of points, the higher heterogeneity of IL-HET can be seen by compaing the means and standard deviations of the loads.
Figure 3: (a) Distribution of building type in IL-HOM. (b) Boxplot of the spread of values in the time series of each building type in IL-HOM.
Further evidence of the homogeneity of IL-HOM can be inferred from the equivalent figures to Figures 2(a) and 2(b) as presented in Figures 3(a) and 3(b). While both figures contain only one element thanks to IL-HOM only containing buildings of the type Warehouse, the lower spread as well as variance of loads is evident from Figure 3(b) as compared to IL-HET.
Table 3: IL-HOM and IL-HET compared to other building energy datasets.
Dataset # buildings Sampling frequency Duration Total load observations Non-time features
IL-HOM 592 15 minutes 1 year 20,755,520 5
IL-HET 592 15 minutes 1 year 20,755,520 5
Electricity Load Diagrams 370 15 minutes 4 years 51,894,720 0
London Smart Meters 5,567 30 minutes $\sim$2 years 168,000,000 0
Buildings-900K 900,000 1 hour 2 years 15,777,000,000 3
We compare IL-HOM and IL-HET to other building-level load datasets, including Electricity Load Diagrams [Trindade, 2015], Low Carbon London [Schofield et al., 2015], and Buildings-900K [Emami et al., 2023]. While other datasets exceed ours in terms of the number of data points, we provide a larger number of features as shown in Table 3, which provides important context to load forecasting models. As a matter of fact, the only other dataset on the list with non-time features is Buildings-900K which includes the latitude/longitude of the building’s PUMA (Public Use Microdata Area) and a binary indicator for whether the building is commercial or residential.
Appendix B Correlation Analysis for Feature Selection
In this section, we discuss our choice of features shown in Section 2. For this analysis, we consider 42 randomly selected Illinois buildings from ComStock which are not in either IL-HOM or IL-HET (which acts as a kind of ‘validation set’ aiding in the construction of the datasets). We split our analysis into two parts: choice of weather (dynamic) features and choice of static features. Weather data available with ComStock includes seven features viz. four temperature/wind features and three heat radiation features. We ignore the latter since they are not available in typical meteorological reports. Among the former four features, we calculate their Pearson correlation coefficient with the load values, and average the correlation across all 42 buildings. The results are presented in Table 4.
Table 4: Correlation analysis of weather features.
Feature name Average correlation coefficient
Dry Bulb Temperature [${}^{\circ}C$] 0.2507
Wind Speed [$m/s$] 0.1396
Wind Direction [Deg] 0.0787
Relative Humidity [%] -0.2794
From Table 4, we select Dry Bulb Temperature [${}^{\circ}C$] and Wind Speed [$m/s$], since Wind Direction [Deg] is almost uncorrelated with load values and Relative Humidity [%] is negatively correlated. On the other hand, ComStock contains many static features to choose from. In order to calculate static features’ correlation to loads, we average the loads across all time indices and calculate the correlation of the averaged loads with the building’s static features.
Table 5: Correlation analysis of static features.
Feature name Correlation coefficient
External window area ($m^{2}$) 0.9714
Floor space ($\text{ft}^{2}$) 0.9245
External wall area ($m^{2}$) 0.7385
Number of air loops 0.7123
Total natural gas consumption 0.7112
Hot water supply volume ($m^{3}$) 0.2672
Occupant equivalent full load hour (EFLH) 0.2672
HVAC cooling type 0.1386
We choose the top three correlated features in order to prevent overcrowding of input data to the models with static features. Thus, the aforementioned analysis informs our choice of features, which in turn determines the models’ learning capabilities.
Appendix C Model Architectures
Table 6: Model sizes, parameter counts and structural hyperparameters for ($T=96,L=48$).
Model name Model hyperparameters Parameter count Model size
LSTM Hidden size: $32$ LSTM Layers: 1 Dropout: 0.1 13,856 0.05 MB
LSTNet Conv. layer output channels: 32 Conv. kernel size: 12 GRU hidden size: 128 Recurrent layer skips: $\{4,4\}$ Attention window size: 7 Dropout: 0.1 66,473 0.25 MB
Transformer Token size: 128 Num. heads: 4 Encoder layers: 3 Decoder layers: 3 FCNN hidden dimension: 512 Activation: GeLU Dropout: 0.1 2,698,625 10.29 MB
Autoformer Token size: 128 Num. heads: 4 Encoder layers: 3 Decoder layers: 3 FCNN hidden dimension: 512 Moving average window: 25 TA Delay top-$k$: $5\log(L)$ Dropout: 0.1 1,413,633 5.39 MB
Informer Token size: 128 Num. heads: 4 Encoder layers: 3 Decoder layers: 3 FCNN hidden dimension: 512 Activation: GeLU Dropout: 0.1 2,798,211 10.67 MB
TimesNet Token size: 128 Encoder layers: 3 $k$ for top-$k$ amplitudes: 5 Num kernels: 3 Inception block intermediate channels: 64 Dropout: 0.1 998,289 3.81 MB
Our choice of base models is meant to capture the major trends towards time series forecasting using deep architectures in recent years. To that end, we summarise these trends and the positioning of our chosen architectures within those trends.
  • Recurrent Neural Network Architectures: These models consist of a learnable kernel that is unrolled recursively in time to generate forecasts. In this category, we use long-short term memory (LSTM) [Hochreiter and Schmidhuber, 1997] with a dense decoder, and LSTNet [Lai et al., 2018], which combines RNNs with convolutional layers.
  • Transformer-based Architectures: These models use self-attention and cross-attention mechanisms in an encoder-decoder framework to parallelize the generation of the output sequence, given an input sequence. In this category, we use the vanilla Transformer [Vaswani et al., 2017] and Informer [Zhou et al., 2021], with the latter incorporating the ProbSparse attention mechanism. Contrary to conventional implementations, we use positional embeddings for temporal inputs (i.e. $\mathbf{u}$), alongside the standard positional embeddings.
  • Decomposition-based Architectures: These models replace the attention mechanism of the Transformer with other mechanisms that leverage inherent characteristics of sequential data. In this category, we use Autoformer [Wu et al., 2021], which uses an auto-correlation mechanism, alongside the decomposition of sequences into seasonal and trend patterns.
  • Patch-based Architectures: These models split the input sequence into smaller sub-sequences called patches, followed by learning the interactions between different patches to better predict the forecast. In this category, we use TimesNet [Wu et al., 2023], which creates 2D frequency-domain patches from 1D input sequences, followed by passing them through convolution-based Inception blocks [Szegedy et al., 2015] to generate the forecast.
A detailed description of the model sizes, parameter counts, and structural hyperparameters for the aforementioned models is provided in Table 6.
TimesFM [Das et al., 2024] is a decoder-only foundation model built for time series forecasting. The model consists of input layers, stacked transformers, and output layers and considers an MSE training loss function. The pre-trained TimesFM model timesfm-1.0-200m is trained on a vast amount and variety of data with both synthetic data and real-world data. There are two major model hyperparameters, namely, model_dim, which determines the size of processed data, and num_layers, which controls the number of stacked transformers in the model. A summary of the model size, parameter counts, and structural hyperparameters in TimesFM is provided in Table 7.
Table 7: Model sizes, parameter counts and structural hyperparameters.
Model name Model hyperparameters Parameter count Model size
TimesFM model_dim: $1280$ num_layers: $20$ 200m 854MB
Appendix D Hyperparameter Tuning
Given these models span a wide range of architectures, using the same value for the learning rate for all of them would lead to poor performance. To that end, we tuned the learning rate over the validation set of the problem $(T=96,L=48)$ with choices $\texttt{lr}=\{10^{-3},5\times 10^{-4},10^{-4},5\times 10^{-5},10^{-5},5\times 10^{-6}\}$. The results are reported in Table 8.
Table 8: Learning rate tuning over the validation set of the problem $(L=96,T=48)$. The metric is MSE, and the best learning rates are highlighted.
Model $10^{-3}$ $5\times 10^{-4}$ $10^{-4}$ $5\times 10^{-5}$ $10^{-5}$ $5\times 10^{-6}$
LSTM 2.3091 5.3035 0.4369 0.3611 0.3807 0.6348
LSTNet 0.9551 0.5223 0.1652 0.1246 0.1573 0.2479
Transformer 9.3107 8.7265 1.5617 0.8678 2.3310 1.2849
Autoformer 0.4911 0.3675 0.2472 0.1634 0.2441 0.3875
Informer 8.7894 10.1551 9.1948 15.2048 2.9153 0.6201
TimesNet 0.0781 0.0845 0.0807 0.0796 0.0773 0.0814
Since Informer showed the best performance at the end of the learning rate choices, we ran a further experiment to test whether even lower learning rates are better for it, whose conclusion was negative (see Table 9).
Table 9: Extended learning rate tuning for Informer.
Model $10^{-6}$ $5\times 10^{-7}$ $10^{-7}$
Informer 2.5689 4.1384 1.2238
The training is carried out using the above model-specific learning rates. As described in Section 3, we use an early stopping criteria of no improvement in the validation set metrics (also called patience) for 5 epochs. On the other hand, all training terminates in 20 epochs whether or not the patience criteria is met. The epoch at which each model’s training terminates is given in Table 10.
Table 10: Termination epoch for each trained model. The asterisk (*) indicates that early stopping was not triggered.
Model IL-HET IL-HOM
$T=4$ $T=48$ $T=96$ $T=4$ $T=48$ $T=96$
LSTM $20^{*}$ 1 $20^{*}$ $20^{*}$ $20^{*}$ $20^{*}$
LSTNet 7 15 5 19 $20^{*}$ $20^{*}$
Transformer 7 2 8 19 $20^{*}$ $20^{*}$
Autoformer 5 1 1 9 15 2
Informer 10 2 2 7 17 15
TimesNet 2 1 19 19 $20^{*}$ $20^{*}$
The observations of Table 10 can be explained by the fact that training of IL-HET is more unstable than IL-HOM, which leads to early stopping being triggered very early into the training run. This is further verified by letting Transformer for $L=96,T=96$ train to completion without early stopping on IL-HET, which reveals instability in training around epoch 8 (see Figure 4).
Figure 4: Training instability in Transformer training on IL-HET.
Similar to the base models, the learning rate for TimeFM was tuned over the validation set of the problem $(T=96,L=96)$ with choices $\texttt{lr}=\{10^{-2},10^{-3},5\times 10^{-4},10^{-4},5\times 10^{-5},10^{-5},5\times 10^{-6}\}$ over 1 epoch. The results are reported in Table 11. Note that $10^{-2}$ is the default learning rate in the pre-trained TimesFM model.
Table 11: Learning rate tuning over the validation set of the problem $(L=96,T=96)$. The metric is MSE, and the best learning rates are highlighted.
Model $10^{-2}$ $10^{-3}$ $5\times 10^{-4}$ $10^{-4}$ $5\times 10^{-5}$ $10^{-5}$ $5\times 10^{-6}$
TimesFM 0.2034 0.4440 0.5235 0.6985 0.7347 0.7566 0.7582
Patience is also implemented in fine-tuning the TimesFM model where validation is performed to compute the progress of the fine-tuning in terms of the MSE loss every 1,000 steps in each epoch. Patience is increased by one if the MSE loss does not decrease and the so-called early termination of the current epoch is executed if patience reaches five. In our fine-tuning experience, the early termination condition was never triggered.
Appendix E Predictions
In Figure 5, we show the predictions for the first 500 points of the $5^{\text{th}}$ customer in both the datasets. These visualizations offer insights into how well the trained models capture the underlying patterns in individual building data. However, it is important to note that the predictions for a single building may not fully represent the model’s performance across the entire test set.
Figure 5: Predictions of the test set loads for two selected clients from IL-HOM and IL-HET datasets, respectively, for the best models reported in Table 6.